# Evaluating Meta-Regression Techniques: A Simulation Study on Heterogeneity in Location and Time

### v1

@@ 4.1 Location-level Effects: Fixed, Random, and Mixed Effects
When key factors of a group of studies vary within each location, a location-level effects model can be used to control the variation between locations. This operates nearly in the same way as the study-level effect model does with the same estimation strategy and underlying assumptions/constraints (please see Section 3 for further details). The only difference is the subscripts that are now at the location-level, $l$, instead of at the study-level, $s$. An important note is that a location-level model should be used instead of a study-level model when there is specific heterogeneity in locations. The above description applies to fixed effects, random effects, or mixed effects; each of these methodology specifications will be reviewed in this section.
Fixed Effects
Equation 5 below represents location-fixed effects.
For location level fixed-effects, each location is indexed by $l=1,...,L$, $y_{l}$ is a $(l\times 1)$ vector for the variable of interest, $\mathbf{x}_{l}^{\top}=(1,x_{l1},x_{l2},...,,x_{lp})$ represents a $(l\times p)$ vector of covariates per location $l$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ represents a $(p\times 1)$ vector of effect sizes (i.e. regression coefficients), and $\varepsilon_{l}$ is a $(l\times 1)$ vector for the error term. Note that $\boldsymbol{\alpha}_{l}$ denotes the fixed effect where $\boldsymbol{\alpha}_{l,FE}=\gamma_{l}D_{l}$ such that $D_{l}$ is a vector of dummy variables per location and $\gamma_{l}$ is a vector of fixed effects for location $l$ (which represents the difference in effect size between each location $l$).
This specification would be most appropriate when the random effects assumption does not hold and heterogeneity is present at only the location level. Note that the variation between locations is omitted, but within-location variation is retained. Also, note that homogeneity in time is still assumed if these location-level fixed effects are not combined with time controls (either a trend or time-level fixed effect). The parameter is estimated using the weights of each group (in this case, each location) as seen in Equation 6 below.
$SE(\hat{\beta})=\sqrt{\frac{1}{\sum^{L}_{l=1}\frac{1}{\sigma^{2}_{l}}}}$ (7)
To estimate the standard error above $SE(\hat{\beta})$, the location-level population variance $\sigma^{2}_{l}$ is needed (Domínguez Islas & Rice, 2018). Grouping by location, $l$, instead of by study, $s$, is a key difference since variation in locations is now explicitly controlled for, instead of being implicitly controlled for by the grouping of studies.
Random Effects
The simulation will also test location-level random effects, which is similar, but has some slight differences. This also uses Equation 5, but the fixed effect is replaced with the random effect $\boldsymbol{\alpha}_{l,RE}$. This random effect is drawn from $(0,\sigma_{\alpha}^{2})$ and the error term $\varepsilon_{l}$ is $(0,\sigma_{\varepsilon}^{2})$ distributed. The estimation of parameters and their standard errors is the same as it was for study-level random effects in Appendix G, just the subscripts change from study-level, $s$ to location-level, $l$ as seen below in Equation 8 & 9. This specification would be most appropriate when the random effects assumption holds, heterogeneity is present at only the location level, and subjects are randomly drawn or represent the full population.
Mixed Effects
In the equation above, $y_{l}$ denotes a $(l\times 1)$ vector for the dependent variable, $\mathbf{x}_{l}^{\top}=(1,x_{l1},x_{l2},...,,x_{lp})$ representing a $(l\times p)$ vector of covariates per location $l$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ representing a $(p\times 1)$ vector of regression coefficients (i.e. effect sizes), $\boldsymbol{\alpha}_{l,FE}$ denotes the fixed effects vector, $\boldsymbol{\alpha}_{l,RE}$ denotes the random effects vector, and $\varepsilon_{l}$ denotes the $(l\times 1)$ error term vector (Viechtbauer, 2007). Estimating the parameters and their standard errors will be the same as it is for location-level random effects for the same reasons articulated for study-level mixed effects in Appendix G (please see Equation 8 & 9). This is because a mixed effect model is comprised of both random effects as well as fixed effects, and fixed effects provide the foundation to calculate random effects.
Summary
In summary, the only difference between the standard methodology used for conducting a meta-analysis and the location-level methodology is the respective use of study-level grouping versus location-level grouping. This holds for fixed, random, or mixed effects. The math notation is very similar for location-level models compared to study-level models, the only difference is the $s$ subscripts are changed with $l$ subscripts. An important note is that a location-level model should be used instead of a study-level model when there is specific heterogeneity in locations.
@@ 4.2 Time-level Fixed Effects
This specification would be most appropriate when heterogeneity is present at only the time level. A limitation to note is that variation across time cannot be accounted for because comparisons are made within each year but not between. Also, if a time-level fixed effect is not combined with a location-level fixed effect, this model then assumes homogeneity in locations, and that homogeneity would need to be tested. Estimation of parameters and their standard errors are the same as it was in Section 4.1 and for study-level fixed effects in Appendix G. The parameter is estimated using the weight of each group (in this case each time) as seen in Equation 12 below.
$SE(\hat{\beta})=\sqrt{\frac{1}{\sum^{T}_{t=1}\frac{1}{\sigma^{2}_{t}}}}$ (13)
To estimate the standard error $SE(\hat{\beta})$, the time-level population variance is needed $\sigma^{2}_{t}$ (Domínguez Islas & Rice, 2018). Grouping by time, $t$ instead of by study, $s$, is a key difference since variation in time is now explicitly controlled for, instead of being implicitly controlled for by the grouping of studies.
Similar to Section 4.1, the only difference between the standard methodology used to conduct a meta-analysis and the time-level methodology is the respective use of study-level grouping versus time-level grouping. This holds for both fixed or random effects, and the only difference in the math notation is the $s$ subscripts are changed with $t$ subscripts. An important note is that a time-level model should be used instead of a study-level model when there is specific heterogeneity in time periods.
@@ 4.3 Combined Location and Time-level Fixed Effects
Note that when both location and time-level fixed effects are combined, multicollinearity needs to be tested since any location may have a direct link to a specific time. For instance, what if there is only one study in the meta-analysis that was conducted in Antarctica? What if this same study was the only one conducted in 2024? In this hypothetical example, the dummy variable for 2024 is perfectly collinear with Antarctica. Also note that this methodology specification has the same underlying assumptions/constraints of study-level fixed effects in Appendix G since it also uses fixed effects. The estimation of parameters and their standard errors will be calculated similarly as in Section 4.1 & 4.2 with the same estimation strategy (see study-level fixed effects in Appendix G for further details). Both indices need to be paired together since each observation is observed at a specific location and time period. The parameter is estimated using the weight of both groupings (location and time) as seen in Equation 15 below.
@@ 4.4 Modeling Trend with Fixed Effects: Study or Location Level
Either methodology specification (study or location level), is achieved by adding a rescaled trend polynomial term $t_{0}=\frac{2t-n-1}{n-1}$ that rescales to $[-1,1]$. The number of time periods is represented by $n$ and each step of a time period is represented by $t=\{1,2,...,\}$. This monotonic transformation is used to combat the near-multicollinearity likely present in trend polynomials, especially when high degrees are included (Spanos & McGuirk, 2002; Spanos, 2019a; Seber & Lee, 2003). Trend polynomials have also been found to be unstable due to an ill-conditioned matrix in estimation unless they are rescaled (Michaelides & Spanos, 2020; Seber & Lee, 2003). This trend, $t_{0}$, is simply one of the covariates, so it is nested inside the same $x_{i}$ for study-level fixed effects in Appendix G and Section 4.4 for location-level fixed effects. To illustrate this trend polynomial transformation suppose your yearly data spans 2020-2024. The values for $t_{0}$ will accordingly be $\{-1,-0.5,0,0.5,1\}$ by using the following values of $t=\{1,2,3,4,5\}$. Now the magnitude is removed so the only factor considered are the time intervals themselves.
4.5 Summary
Overall, the beginning of Section 4 introduced specifications that control for location and/or time heterogeneity. The subsections showed the details of each specification including location-level fixed/random/mixed effects, time-level fixed effects, combined location and time level fixed effects, and location-level fixed effects with a trend. These methodology specifications are similar in that they control for heterogeneity in location and/or time, but the difference lies in how each of them are done. Note that location-level fixed effects has been proposed in the experimental economics literature, but this methods performance has not been fully assessed in a simulation until now. Also, although time-level fixed effects have been tested in the meta-analysis simulation literature, study-level or location-level fixed effects with a trend has not been tested until now. The following section will show details of the simulation including its design, how its performance is assessed, and its results.
@@ 5.1 Design
The standard Monte Carlo simulation is conducted with 10,000 iterations. For each Monte Carlo iteration, the meta-regression data has the following parameters as specified in Table 1: the sample size in each study (50, 100 & 150), the number of covariates (i.e. independent variables) in each study (1, 3 & 5), the number of locations (5, 9, 15), and the number of studies (25, 45, 75) in the meta-regression. Each location in the meta-regression include 1 study per year for 5 years, which results in either 25, 45, or 75 studies due to the number of locations (5, 9, or 15). Excluding the number of locations, these are common parameters tested in the literature with values varying depending on the research question or the measure that is being tested (Richardson et al., 2021; Sugasawa & Noma, 2021; Dormann et al., 2020; Field, 2001; Kulinskaya et al., 2021; Bakbergenuly and Kulinskaya, 2018). Note that the meta-regression data does not have any overlap across studies, which means that no data from the same study is double counted in the meta-regression. This is the most common case when conducting a meta-regression (Lin & Sullivan, 2009; Bom & Rachinger, 2020; Hennessy & Johnson, 2020). Practitioners need to check if there is overlap or not though because if there is, this could introduce confounding factors in the model because of a duplicate study in the meta-regression. Also note that no data points are missing in the meta-regression data. This is unlikely in practice and will be explored in future research. Each parameter always has a range of values to test a variety of situations, and the ranges of values we selected are very commonly used in the literature. The purpose of these parameters is to capture if variation in each of these parameters impacts the performance of each method.
@@
Parameter Symbol Values
Sample size (per study in meta-regression) $n$ 50, 100, 150
Number of covariates $k$ 1, 3, 5
Number of locations (in meta-regression) $n_{l}$ 5, 9, 15
Number of studies (in meta-regression) $n_{s}$ 25, 45, 75
@@
To simulate location and time level heterogeneity, our simulation includes twelve cases that embed varying degrees of location and time level heterogeneity into the mean $\mu_{Y}$ when generating the simulated data for the dependent variable $Y$. Table 2 shows the differences between all cases, which are as follows. Cases 1-6 include a small time effect by adding 0.1 to $\mu_{Y}$ each year, while Cases 7-12 include a large time effect by adding 0.5 to $\mu_{Y}$ each year. This means that 0.1 (or 0.5) is added once to $\mu_{Y}$ in the second year, but added a total of four times in the fifth year, resulting in 0.4 (or 2.0) added to $\mu_{Y}$ over the five years. Cases 1-2 & 7-8 include 5 locations, Cases 3-4 & 9-10 include 9 locations, while the remaining cases include 15 locations. The odd numbered cases include a small location effect by assigning each location with a different $\mu_{Y}$; 5 locations are assigned one element each from $\mu_{Y}=\{-2,-1,0,1,2\}$, 9 locations are assigned one element each from $\mu_{Y}=\{-2,-1.5,-1,-0.5,0,0.5,1,1.5,2\}$, and 15 locations are assigned one element from $\mu_{Y}=\{0,\pm(0.15,0.33,0.66,1,1.33,1.66,2)\}$. In contrast, the even-numbered cases include a large location effect by assigning each location with a different $\mu_{Y}$; 5 locations are assigned one element from $\mu_{Y}=\{-10,-5,0,5,10\}$, 9 locations are assigned one element from $\mu_{Y}=\{-10,-7.5,-5,-2.5,0,2.5,5,7.5,10\}$, and 15 locations are assigned one element from $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10)\}$. Note that the $\mu_{Y}$ values are not based on real data, they are solely illustrative of small versus large differences in the means of the dependent variable in each location. The large versus small differences in the mean of each location simulate large versus small effects of location-level heterogeneity. Similarly, the smaller versus larger trend added to the mean simulates weaker versus more powerful effects of time level heterogeneity. Note that both trends are designed to be positive, constant over time, and at the same rate for all locations. These differences in generating the dependent variable test how well each of the methodology specifications performs in the face of varying degrees of both location and time level heterogeneity. Section 5.2 defines specific model performance criteria.
Case Locations Time Effect Location Effect
1 5 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-2,-1,0,1,2\}$
2 5 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-10,-5,0,5,10\}$
3 9 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{-2,-1.5,-1,0.5,0,0.5,1,1.5,2\}$
6 15 $\mu_{Y}=\mu_{Y}+0.1\cdot t$ $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10))\}$
7 5 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-2,-1,0,1,2\}$
8 5 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-10,-5,0,5,10\}$
9 9 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{-2,-1.5,-1,0.5,0,0.5,1,1.5,2\}$
12 15 $\mu_{Y}=\mu_{Y}+0.5\cdot t$ $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10)\}$
@@
When generating the covariates for each of the studies in the meta-regression, we generate them as Normal, Independent, and Identically-Distributed with a pre-specified variance-covariance matrix. By generating data with a specified variance-covariance matrix, the true parameter estimates can be calculated (which is needed to test the accuracy of each model’s predictions), and our covariates will be statistically significant ensuring that our estimators have enough power (Spanos & McGuirk, 2001; Sugasawa & Noma, 2021; Dormann et al., 2020; Field, 2001). Note that the Normal distribution was chosen since most software packages analyze random effects and fixed effects models by assuming the error term or random effect are normally distributed to be able to hypothesis test and generate confidence intervals via the central limit theorem (Pateras et al., 2018; Pereira et al., 2010; Sánchez-Meca & Marín-Martínez, 2008; Cameron & Trivedi, 2005; Wooldridge, 2010). The literature has demonstrated that estimators that assume a normal distribution (i.e. majority of estimators) perform very poorly when the data are non-normally distributed, even when the sample size is large (Blázquez-Rincón et al., 2023; McGrath et al., 2023; Bohnet, 2007; Bergtold et al., 2010; Poudyal & Spanos, 2022; Rubio‐Aparicio et al., 2018). The mean of each covariate is assigned a value close to the range of $[0,1]$ following the suggested design from the literature (Richardson et al., 2021; Blázquez-Rincón et al., 2023; Pateras et al., 2018; McGrath et al., 2019; Sánchez-Meca & Marín-Martínez, 2008). Variances of 1 are also suggested by the literature for simplicity. These parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from said data. Most studies extract parameters from one particular meta-regression to ensure the parameters are consistent as much as possible (Pereira et al., 2010; Schwarzer et al., 2024; Dormann et al., 2020). To analyze beyond the sample, one study extracted parameters from 12,894 meta-analyses in the Cochrane Database of Systematic Reviews (Langan et al., 2019). Figure 3 in Appendix B presents the summary statistics for all variables (five covariates) as box and whisker plots. This is done to help visualize the data generation process.
Since $\mu_{Y}$ varies per study, each study in the meta-regression has a different true intercept coefficient, $\beta_{0}$, calculated by Equation 17 below.
@@
To recap the design of this standard Monte Carlo simulation, Figure 1 shows detailed steps of how the simulation was conducted starting from data generation and ending in how the results are evaluated with the performance criteria. Table 1 shows the parameters used when generating the meta-regression data, and Table 2 shows how small or large heterogeneity in location or time are generated through the mean of the variable of interest ($\mu_{Y}$) when generating the data. when generating the meta-regression data we generate our variables using a Normal, Independent and Identically-Distributed joint distribution with a pre-specified variance-covariance. This allows us to calculate the true parameter value which is needed to test the accuracy of each model’s predictions, and since our covariates will be statistically significant our estimators will have appropriate power. The mean and variance values are respectively within $[0.1]$ and 1 as suggested by the literature, but note that these mean and variance parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from the real data sets (Richardson et al., 2021; Blázquez-Rincón et al., 2023; Pateras et al., 2018; McGrath et al., 2019; Sánchez-Meca & Marín-Martínez, 2008). The following subsections show in detail how the true value of the coefficients are calculated for when there are one, three, or five covariates.

### v2

@@ 4.1 Location-level Effects: Fixed, Random, and Mixed Effects
When key factors of a group of studies vary within each location, a location-level effects model can be used to control the variation between locations. This operates nearly in the same way as the study-level effect model does with the same estimation strategy and underlying assumptions/constraints (please see Section 3 for further details). The only difference is the subscripts that are now at the location-level, $l$, instead of at the study-level, $s$. An important note is that a location-level model should be used instead of a study-level model when there is specific heterogeneity in locations. The above description applies to fixed effects, random effects, or mixed effects. Note that random effects and mixed effects specifications are only appropriate when the random effects assumption holds and subjects are randomly drawn or represent the full population. In contrast, only fixed effects would be appropriate if the random effects assumption does not hold. Note that the variation between locations is omitted, but within-location variation is retained. Also, note that homogeneity in time is still assumed if these location-level fixed effects are not combined with time controls (either a trend or time-level fixed effect).
@@ 4.2 Time-level Fixed Effects
This specification would be most appropriate when heterogeneity is present at only the time level. A limitation to note is that variation across time cannot be accounted for because comparisons are made within each year but not between. Also, if a time-level fixed effect is not combined with a location-level fixed effect, this model then assumes homogeneity in locations, and that homogeneity would need to be tested. Similar to Section 4.1, the only difference between the standard methodology used to conduct a meta-analysis and the time-level methodology is the respective use of study-level grouping versus time-level grouping. This holds for both fixed or random effects, and the only difference in the math notation is the $s$ subscripts are changed with $t$ subscripts. An important note is that a time-level model should be used instead of a study-level model when there is specific heterogeneity in time periods.
@@ 4.3 Combined Location and Time-level Fixed Effects
Note that when both location and time-level fixed effects are combined, multicollinearity needs to be tested since any location may have a direct link to a specific time. For instance, what if there is only one study in the meta-analysis that was conducted in Antarctica? What if this same study was the only one conducted in 2024? In this hypothetical example, the dummy variable for 2024 is perfectly collinear with Antarctica. Also note that this methodology specification has the same underlying assumptions/constraints of study-level fixed effects in Appendix G since it also uses fixed effects.
@@ 4.4 Modeling Trend with Fixed Effects: Study or Location Level
Either methodology specification (study or location level), is achieved by adding a rescaled trend polynomial term $t_{0}=\frac{2t-n-1}{n-1}$ that rescales to $[-1,1]$. The number of time periods is represented by $n$ and each step of a time period is represented by $t=\{1,2,...,\}$. This monotonic transformation is used to combat the near-multicollinearity likely present in trend polynomials, especially when high degrees are included (Spanos, 2019a; Seber & Lee, 2003; Spanos & McGuirk, 2002). Trend polynomials have also been found to be unstable due to an ill-conditioned matrix in estimation unless they are rescaled (Michaelides & Spanos, 2020; Seber & Lee, 2003). This trend, $t_{0}$, is simply one of the covariates, so it is nested inside the same $x_{i}$ for study-level fixed effects in Appendix G and Section 4.4 for location-level fixed effects. To illustrate this trend polynomial transformation suppose your yearly data spans 2020-2024. The values for $t_{0}$ will accordingly be $\{-1,-0.5,0,0.5,1\}$ by using the following values of $t=\{1,2,3,4,5\}$. Now the magnitude is removed so the only factor considered are the time intervals themselves.
@@ 5.1 Design
The standard Monte Carlo simulation is conducted with 10,000 iterations. For each Monte Carlo iteration, the meta-regression data has the following parameters as specified in Table 1: the sample size in each study (50, 100 & 150), the number of covariates (i.e. independent variables) in each study (1, 3 & 5), the number of locations (5, 9, 15), and the number of studies (25, 45, 75) in the meta-regression. Each location in the meta-regression include 1 study per year for 5 years, which results in either 25, 45, or 75 studies due to the number of locations (5, 9, or 15). Excluding the number of locations, these are common parameters tested in the literature with values varying depending on the research question or the measure that is being tested (Richardson et al., 2021; Sugasawa & Noma, 2021; Dormann et al., 2020; Kulinskaya et al., 2021; Bakbergenuly and Kulinskaya, 2018; Field, 2001). Note that the meta-regression data does not have any overlap across studies, which means that no data from the same study is double counted in the meta-regression. This is the most common case when conducting a meta-regression (Bom & Rachinger, 2020; Hennessy & Johnson, 2020; Lin & Sullivan, 2009). Practitioners need to check if there is overlap or not though because if there is, this could introduce confounding factors in the model because of a duplicate study in the meta-regression. Also note that no data points are missing in the meta-regression data. This is unlikely in practice and will be explored in future research. To better visualize this type of dataset, suppose there are five time periods for five locations, say 2020-2024. This would mean that there are $25$ studies total in the meta-regression since $L=5$ and $T=5$. Since each study includes subject-level data, the compiled meta-regression data set has sample size $N=25n$ where n is the number of subjects in the study. For instance, if $n=100$, $N=2,500$.
@@
Each parameter always has a range of values to test a variety of situations, and the ranges of values we selected are very commonly used in the literature. The purpose of these parameters is to capture if variation in each of these parameters impacts the performance of each method. To simulate location and time level heterogeneity, our simulation includes twelve cases that embed varying degrees of location and time level heterogeneity into the mean $\mu_{Y}$ when generating the simulated data for the dependent variable $Y$. Table 2 shows the differences between all cases, which are as follows. Cases 1-6 include a small time effect by adding 0.1 to $\mu_{Y}$ each year, while Cases 7-12 include a large time effect by adding 0.5 to $\mu_{Y}$ each year. This means that 0.1 (or 0.5) is added once to $\mu_{Y}$ in the second year, but added a total of four times in the fifth year, resulting in 0.4 (or 2.0) added to $\mu_{Y}$ over the five years. Cases 1-2 & 7-8 include 5 locations, Cases 3-4 & 9-10 include 9 locations, while the remaining cases include 15 locations. The odd numbered cases include a small location effect by assigning each location with a different $\mu_{Y}$; 5 locations are assigned one element each from $\mu_{Y}=\{-2,-1,0,1,2\}$, 9 locations are assigned one element each from $\mu_{Y}=\{-2,-1.5,-1,-0.5,0,0.5,1,1.5,2\}$, and 15 locations are assigned one element from $\mu_{Y}=\{0,\pm(0.15,0.33,0.66,1,1.33,1.66,2)\}$. In contrast, the even-numbered cases include a large location effect by assigning each location with a different $\mu_{Y}$; 5 locations are assigned one element from $\mu_{Y}=\{-10,-5,0,5,10\}$, 9 locations are assigned one element from $\mu_{Y}=\{-10,-7.5,-5,-2.5,0,2.5,5,7.5,10\}$, and 15 locations are assigned one element from $\mu_{Y}=\{0,\pm(1,2.5,4,5.5,7,8.5,10)\}$. Note that the $\mu_{Y}$ values are not based on real data, they are solely illustrative of small versus large differences in the means of the dependent variable in each location. The large versus small differences in the mean of each location simulate large versus small effects of location-level heterogeneity. Similarly, the smaller versus larger trend added to the mean simulates weaker versus more powerful effects of time level heterogeneity. Note that both trends are designed to be positive, constant over time, and at the same rate for all locations. These differences in generating the dependent variable test how well each of the methodology specifications performs in the face of varying degrees of both location and time level heterogeneity. Section 5.2 defines specific model performance criteria.
@@
When generating the covariates for each of the studies in the meta-regression, we generate them as Normal, Independent, and Identically-Distributed with a pre-specified variance-covariance matrix. By generating data with a specified variance-covariance matrix, the true parameter estimates can be calculated (which is needed to test the accuracy of each model’s predictions), and our covariates will be statistically significant ensuring that our estimators have enough power (Sugasawa & Noma, 2021; Dormann et al., 2020; Spanos & McGuirk, 2001; Field, 2001). Note that the Normal distribution was chosen since most software packages analyze random effects and fixed effects models by assuming the error term or random effect are normally distributed to be able to hypothesis test and generate confidence intervals via the central limit theorem (Pateras et al., 2018; Pereira et al., 2010; Wooldridge, 2010; Sánchez-Meca & Marín-Martínez, 2008; Cameron & Trivedi, 2005). The literature has demonstrated that estimators that assume a normal distribution (i.e. majority of estimators) perform very poorly when the data are non-normally distributed, even when the sample size is large (Blázquez-Rincón et al., 2023; McGrath et al., 2023; Poudyal & Spanos, 2022; Rubio‐Aparicio et al., 2018; Bergtold et al., 2010; Bohnet, 2007). The mean of each covariate is assigned a value close to the range of $[0,1]$ following the suggested design from the literature (Blázquez-Rincón et al., 2023; Richardson et al., 2021; McGrath et al., 2019; Pateras et al., 2018; Sánchez-Meca & Marín-Martínez, 2008). Variances of 1 are also suggested by the literature for simplicity. These parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from said data. Most studies extract parameters from one particular meta-regression to ensure the parameters are consistent as much as possible (Schwarzer et al., 2024; Dormann et al., 2020; Pereira et al., 2010). To analyze beyond the sample, one study extracted parameters from 12,894 meta-analyses in the Cochrane Database of Systematic Reviews (Langan et al., 2019). Figure 3 in Appendix B presents the summary statistics for all variables (five covariates) as box and whisker plots. This is done to visualize the data generation process. Since $\mu_{Y}$ varies per study, each study in the meta-regression has a different true intercept coefficient, $\beta_{0}$, calculated by Equation 7 below.
@@
To recap the design of this standard Monte Carlo simulation, Figure 1 shows detailed steps of how the simulation was conducted starting from data generation and ending in how the results are evaluated with the performance criteria. Table 1 shows the parameters used when generating the meta-regression data, and Table 2 shows how small or large heterogeneity in location or time are generated through the mean of the variable of interest ($\mu_{Y}$) when generating the data. when generating the meta-regression data we generate our variables using a Normal, Independent and Identically-Distributed joint distribution with a pre-specified variance-covariance. This allows us to calculate the true parameter value which is needed to test the accuracy of each model’s predictions, and since our covariates will be statistically significant our estimators will have appropriate power. The mean and variance values are respectively within $[0.1]$ and 1 as suggested by the literature, but note that these mean and variance parameters can be linked to real meta-regression data sets if they are substituted with extracted parameters from the real data sets (Blázquez-Rincón et al., 2023; Richardson et al., 2021; McGrath et al., 2019; Pateras et al., 2018; Sánchez-Meca & Marín-Martínez, 2008). The following subsections show in detail how the true value of the coefficients are calculated for when there are one, three, or five covariates.